OpenStudy (kainui):

Here's an interesting equation for complex numbers I came across while playing around:

sammixboo (sammixboo):

Oh boy

OpenStudy (kainui):

\[\Huge z^*=iz\] Solve for z.

Nnesha (nnesha):

* what is this ??????

OpenStudy (unklerhaukus):

complex conjugate

OpenStudy (unklerhaukus):

Assume the complex number \(z\) is of the form \[z=a+ib\] The complex conjugate is \[z^\ast=a-ib\] Plugging into the formula\[a-ib=i(a+ib)\]\[a-ib=-b+ia\] comparing real and complex coefficients\[a=-b\] Hence \[z=a-ia=a(1-i)\] where \(a\) is any real number.

OpenStudy (fibonaccichick666):

but... \[x+iy=i(x-iy)\] isn't true

OpenStudy (fibonaccichick666):

unless what unkle said is true

OpenStudy (kainui):

See, this is what makes it quite interesting. Here is how I originally got here: \[\LARGE z=\sqrt{i}\] Now, \[\LARGE z^* = \sqrt{-i}=\sqrt{-1}\sqrt{i}=i \sqrt{i}=i z\] It is slightly easier to look at @FibonacciChick666 if you consider \[\LARGE z=re^{i \theta}\] rather than a+bi.

OpenStudy (fibonaccichick666):

still, you are assuming a z value

OpenStudy (kainui):

No, I was just showing a particular value of z that worked, but in general there are infinite solutions along the line y=x in the complex plane.

OpenStudy (fibonaccichick666):

but that is not every z, if it doesn't work for every z you have to put conditions on it

OpenStudy (kainui):

It definitely appears that in order to be true like how @UnkleRhaukus solved it that we'll get this formula: \[\LARGE a+b=(a+b)*i\] so that it seems to be implying \[\LARGE 1=i\]

ganeshie8 (ganeshie8):

\[\large e^{-i\theta } = e^{i(\pi/2+\theta) }\]

OpenStudy (fibonaccichick666):

but that isn't the question, it is (a-ib)i

OpenStudy (kainui):

There is no difference, just your choice of coordinates. \[\LARGE z= a+bi = re^{i \theta}\]

OpenStudy (fibonaccichick666):

your secondary, polar form exists for every number

OpenStudy (kainui):

Show us how, I don't believe you. =)

OpenStudy (fibonaccichick666):

no way am I replicating that proof again. You can google it haha I believe euler developed it?

OpenStudy (kainui):

Prove what? I don't understand what you're talking about, I don't think you understand.

OpenStudy (fibonaccichick666):

"Show us how, I don't believe you. =)" I was referring to that.

OpenStudy (anonymous):

\[z \times z^*\] what if you had this

OpenStudy (anonymous):

\[\sqrt{i} \times \sqrt{-i}\]

OpenStudy (kainui):

Yes so does a+bi. It too exists for every number @FibonacciChick666

ganeshie8 (ganeshie8):

\[z\times z^* = |z|\]

OpenStudy (anonymous):

^2?

OpenStudy (anonymous):

Wait couldn't you still use the complex conjugate method

ganeshie8 (ganeshie8):

yeah put a square

OpenStudy (fibonaccichick666):

Let me further extrapolate, I am saying that I am not deriving polar form of a complex number because you said that you do not buy it because it is readily available online. I am not saying they don't exist, I just mean that in this case, if the formula does not work for every value of z, you must put conditions on it or a simple counter example is all we need to disprove the theory

OpenStudy (kainui):

 These are equivalent \[\LARGE re^{i \theta} = r \cos \theta + i r \sin \theta = a+i b\] and I'm just using these facts:\[\LARGE a= r \cos \theta \\ \LARGE b= r \sin \theta\]

ganeshie8 (ganeshie8):

z=iz* is an equation which need not be true for all z right ? polar/rectangular are just two equivalent representations of complex space

OpenStudy (kainui):

Exactly, it's an equation not an identity.

ganeshie8 (ganeshie8):

i think the other equality `zz* = |z|^2` is an identity

OpenStudy (kainui):

Yeah that is haha. Does that clear it up @FibonacciChick666 ?

OpenStudy (kainui):

@UnkleRhaukus why does this equation have a solution when you rearrange this though? \[\LARGE a-ib = -b+ia\] to \[\LARGE (a+b)=i(a+b)\]

OpenStudy (fibonaccichick666):

that wasn't my point, but anyways, if you wish to use an equation, unless you are solving for all possible solutions then you need limitations because it is not always true

OpenStudy (unklerhaukus):

b=-a

OpenStudy (kainui):

I see, because we actually are dividing by zero here if we divide both sides by a+b haha.

OpenStudy (unklerhaukus):

go directly to jail,

OpenStudy (kainui):

IT'S JUST TOO TEMPTING!!! lol

OpenStudy (kainui):

I was just surprised when taking the complex conjugate of sqrt(i) it was equivalent to multiplying by i, which just sort of felt weird to me and kind of reminds me of eigenfunctions.

OpenStudy (kainui):

I suppose in general, \[\LARGE z^* = e^{i \phi}z\] the answer is \[\LARGE z=e^{-i \frac{\phi}{2}}\] to within a scalar multiple. Hmm ok I guess I'll go back to playing around now lol

OpenStudy (ikram002p):

nice